# Optimal estimation for regression discontinuity design

### v1

@@ Appendix A Proofs
$\leq$ $C^{2}\left(\frac{1}{N(\epsilon)}\sum_{i=1}^{N(\epsilon)}\|R_{i}\|\right)^{2}+\frac{1}{4N(\epsilon)}$
$\leq$ $C^{2}\epsilon^{2}+\frac{1}{4N(\epsilon)}\ \leq\ C^{2}\epsilon^{2}+\frac{1}{4(c_{0}n\epsilon-n^{1-\alpha})},$
@@
and the second inequality holds since $\|R_{i}\|\leq\epsilon$ for $i=1,\dots,N(\epsilon)$. If we set $\epsilon=O(n^{-1/3})$ satisfying $\epsilon\geq 2c_{0}^{-1}n^{-\alpha}$, which exists for $\alpha\geq 1/3$, then the right-hand side becomes $O(n^{-2/3})$. For example, if we set $\epsilon=2c_{0}^{-1}n^{-1/3}$, which satisfies $\epsilon\geq 2c_{0}^{-1}n^{-\alpha}$ for $\alpha\geq 1/3$, then the right-hand side becomes
$4C^{2}c_{0}^{-2}n^{-2/3}+\frac{1}{4(2n^{2/3}-n^{1-\alpha})}=\left(4C^{2}c_{0}^{-2}+\frac{1}{4(2-n^{1/3-\alpha})}\right)n^{-2/3}=O(n^{-2/3}).$
Hence, $\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})=O(n^{-2/3})$.
@@
where the last inequality follows from $\theta_{i}+1/2\geq 0$. This implies that $\hat{u}$ converges to one because $\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})$ converges to zero. From Lemma 3.1, we obtain
[…]
Hence, it suffices to show that
$\frac{C^{2}\sum_{i=1}^{n}\hat{w}_{i}^{2}\|R_{i}\|^{2}}{\frac{1}{4}\sum_{i=1}^{n}\hat{w}_{i}^{2}}\ \to\ 0.$ (A.1)
@@
$\frac{C^{2}\sum_{i=1}^{n}\hat{w}_{i}^{2}\|R_{i}\|^{2}}{\frac{1}{4}\sum_{i=1}^{n}\hat{w}_{i}^{2}}$ $=$ $\frac{4C^{2}\sum_{i=1}^{N(\epsilon)}\hat{w}_{i}^{2}\|R_{i}\|^{2}+4C^{2}\sum_{i=N(\epsilon)+1}^{n}\hat{w}_{i}^{2}\|R_{i}\|^{2}}{\sum_{i=1}^{n}\hat{w}_{i}^{2}}$ (A.2)
$\leq$ $4C^{2}\left\{\frac{\epsilon^{2}\left(\sum_{i=1}^{N(\epsilon)}\hat{w}_{i}^{2}\right)+\hat{w}_{N(\epsilon)}^{2}\left(\sum_{i=N(\epsilon)+1}^{n}\|R_{i}\|^{2}\right)}{\sum_{i=1}^{n}\hat{w}_{i}^{2}}\right\}$
$\leq$ $4C^{2}\left\{\epsilon^{2}+\frac{n\hat{w}_{N(\epsilon)}^{2}}{\sum_{i=1}^{n}\hat{w}_{i}^{2}}\right\},$
where the first inequality follows since $\|R_{i}\|\leq\epsilon$ for $i=1,\dots,N(\epsilon)$ and $\hat{w}_{i}\leq\hat{w}_{N(\epsilon)}$ for $i=N(\epsilon)+1,\ldots,n$, and the second inequality follows since $\|R_{1}\|\leq\ldots\leq\|R_{n}\|\leq 1$.
@@
$O(n^{-2/3})$ $=$ $\max_{\bm{\theta}\in\Theta}\text{MSE}(\hat{\bm{w}},\bm{\theta})\ \geq\ \text{MSE}(\hat{\bm{w}},\tilde{\bm{\theta}}(0))$
$\geq$ $C^{2}\left(\sum_{i=1}^{n}\hat{w}_{i}\|R_{i}\|\right)^{2}\ \geq\ C^{2}\hat{w}_{N(\epsilon)}^{2}\left(\sum_{i=1}^{N(\epsilon)}\|R_{i}\|\right)^{2},$
where the second and third inequalities follow from Lemmas 2.3 and 2.2, respectively. Below, we show that, if $\epsilon>2c_{0}^{-1}n^{-\alpha}$, then $\sum_{i=1}^{N(\epsilon)}\|R_{i}\|\geq\frac{(N(\epsilon)-n^{1-\alpha})^{2}}{2c_{1}n}$. Once it is shown, it follows from the assumption $N(\epsilon)\geq c_{0}n\epsilon-n^{1-\alpha}$ that
$O(n^{-2/3})$ $\geq$ $C^{2}\hat{w}_{N(\epsilon)}^{2}\frac{(N(\epsilon)-n^{1-\alpha})^{4}}{4c_{1}^{2}n^{2}}$
@@
$\hat{w}_{N(\epsilon)}^{2}\leq\frac{c_{2}n^{-2/3}}{(c_{0}n^{1/2}\epsilon-2n^{1/2-\alpha})^{4}}\ \ \text{if }\epsilon>2c_{0}^{-1}n^{-\alpha}.$
Now we show the aforementioned claim: if $\epsilon>2c_{0}^{-1}n^{-\alpha}$, $\sum_{i=1}^{N(\epsilon)}\|R_{i}\|\geq\frac{(N(\epsilon)-n^{1-\alpha})^{2}}{2c_{1}n}$. Let $x^{*}(\epsilon)\equiv\frac{N(\epsilon)-n^{1-\alpha}}{c_{1}n}$, which is the unique solution to $c_{1}nx+n^{1-\alpha}=N(\epsilon)$ with respect to $x$. If $\epsilon>2c_{0}^{-1}n^{-\alpha}$, we must have $0<x^{*}(\epsilon)\leq\|R_{N(\epsilon)}\|$, since $c_{1}n\cdot 0+n^{1-\alpha}<c_{0}n\epsilon-n^{1-\alpha}\leq N(\epsilon)$ and $c_{1}n\|R_{N(\epsilon)}\|+n^{1-\alpha}\geq N(\|R_{N(\epsilon)}\|)=N(\epsilon)$ under Assumption 3.1. Here, $N(\|R_{N(\epsilon)}\|)=N(\epsilon)$ holds by the definition of $N(\epsilon)$. Also, let
@@
Note that $g_{\epsilon}(x)\geq N(x)$ for all $x\in(0,\|R_{N(\epsilon)}\|]$, since $g_{\epsilon}(x)=c_{1}nx+n^{1-\alpha}\geq N(x)$ if $0<x\leq x^{*}(\epsilon)$ by Assumption 3.1 and $g_{\epsilon}(x)=N(\epsilon)=N(\|R_{N(\epsilon)}\|)\geq N(x)$ if $x^{*}(\epsilon)<x\leq\|R_{N(\epsilon)}\|$. Therefore, we have
@@
Finally, combining the lower bound on $\sum_{i=1}^{n}\hat{w}_{i}^{2}$ and the upper bound on $\hat{w}_{N(\epsilon)}^{2}$ obtained above yields the following bound on the right-hand side of (A.2): if $\epsilon>2c_{0}^{-1}n^{-\alpha}$,
$4C^{2}\left\{\epsilon^{2}+\frac{n\hat{w}_{N(\epsilon)}^{2}}{\sum_{i=1}^{n}\hat{w}_{i}^{2}}\right\}$ $\leq$ $4C^{2}\left\{\epsilon^{2}+\frac{c_{2}n^{4/3}}{(c_{0}n^{1/2}\epsilon-2n^{1/2-\alpha})^{4}}\hat{u}^{-2}\right\}$
$=$ $4C^{2}\left\{\epsilon^{2}+\frac{c_{2}}{(c_{0}n^{1/6}\epsilon-2n^{1/6-\alpha})^{4}}\hat{u}^{-2}\right\}.$
If we set $\epsilon=n^{-\beta}$ for some $0<\beta<\min\{1/6,\alpha\}$ (which is equivalent to $0<\beta<1/6$ as $\alpha=1/3$), then $\epsilon>2c_{0}^{-1}n^{-\alpha}$ for any sufficiently large $n$, and we have
$4C^{2}\left\{\epsilon^{2}+\frac{c_{2}}{(c_{0}n^{1/6}\epsilon-2n^{1/6-\alpha})^{4}}\hat{u}^{-2}\right\}=4C^{2}\left\{n^{-2\beta}+\frac{c_{2}}{(c_{0}n^{1/6-\beta}-2n^{1/6-\alpha})^{4}}\hat{u}^{-2}\right\}=o(1).$
@@
where $\bm{Y}$ and $\tilde{\bm{Y}}$ follow Bernoulli distribution with parameters $\bm{p}$ and $\tilde{\bm{p}}(p_{0,+},\tau_{0})$, respectively. Then, $\hat{\tau}(\bm{Y})$ is increasing function in $Y_{i,+}$ and decreasing in $Y_{i,-}$. Because $\tilde{Y}_{i,+}$ has first-order stochastic dominance over $Y_{i,+}$ and $\tilde{Y}_{i,-}$ has first-order stochastic dominance over $Y_{i,-}$, it follows from Lemma 1 of Ishihara (2023) that we have
@@ Appendix B Minimax estimation for the average treatment effect
Fix $\theta_{0,+}\in\Theta_{+}$ and $\theta_{0,-}\in\Theta_{-}$. Without loss of generality, we assume $\theta_{0,+}\leq 0$. First, we consider the case where $\theta_{0,+}\in[-1/2,0]$ and $\theta_{0,-}\in[0,1/2]$. In this case, Theorem 2.1 implies that $b_{+}(\bm{w}_{+},\bm{\theta}_{+})^{2}+V_{+}(\bm{w}_{+},\bm{\theta}_{+})$ is maximized at $\bm{\theta}_{+}=\tilde{\bm{\theta}}_{+}(\theta_{0,+})$ and $b_{-}(\bm{w}_{-},\bm{\theta}_{-})^{2}+V_{-}(\bm{w}_{-},\bm{\theta}_{-})$ is maximized at $\bm{\theta}_{-}=\tilde{\bm{\theta}}_{-}(\theta_{0,-})$. Letting
$\bar{\bm{\theta}}_{+}(\theta_{0,+})$ $\equiv$ $\left(\theta_{0,+},\max\{\theta_{0,+}-C\|R_{1,+}\|,1/2\},\ldots,\max\{\theta_{0,+}-C\|R_{n_{+},+}\|,1/2\}\right)^{\prime},$
$\bar{\bm{\theta}}_{-}(\theta_{0,-})$ $\equiv$ $\left(\theta_{0,-},\min\{\theta_{0,-}+C\|R_{1,-}\|,-1/2\},\ldots,\min\{\theta_{0,-}+C\|R_{n_{-},-}\|,-1/2\}\right)^{\prime},$
then $|b_{+}(\bm{w}_{+},\bm{\theta}_{+})|$ is maximized at $\bm{\theta}_{+}=\tilde{\bm{\theta}}_{+}(\theta_{0,+})$ or $\bar{\bm{\theta}}_{+}(\theta_{0,+})$ and $|b_{-}(\bm{w}_{-},\bm{\theta}_{-})|$ is maximized at $\bm{\theta}_{-}=\tilde{\bm{\theta}}_{-}(\theta_{0,-})$ or $\bar{\bm{\theta}}_{-}(\theta_{0,-})$. Because $\bm{w}_{+}\in\mathcal{W}_{+}$ and $\theta_{0,+}\leq 0$, we have $|b_{+}(\bm{w}_{+},\bar{\bm{\theta}}_{+}(\theta_{0,+}))|\leq|b_{+}(\bm{w}_{+},\tilde{\bm{\theta}}_{+}(\theta_{0,+}))|$. Similarly, we have $|b_{-}(\bm{w}_{-},\bar{\bm{\theta}}_{-}(\theta_{0,-}))|\leq|b_{-}(\bm{w}_{-},\tilde{\bm{\theta}}_{-}(\theta_{0,-}))|$. These results imply that $b_{+}(\bm{w}_{+},\bm{\theta}_{+})b_{-}(\bm{w}_{-},\bm{\theta}_{-})$ is maximized at $\bm{\theta}_{+}=\tilde{\bm{\theta}}_{+}(\theta_{0,+})$ and $\bm{\theta}_{-}=\tilde{\bm{\theta}}_{-}(\theta_{0,-})$. Therefore, if $\theta_{0,+}\in[-1/2,0]$ and $\theta_{0,-}\in[0,1/2]$ hold, we obtain
@@
Because $b_{-}(\bm{w}_{-},\bm{\theta}_{-})=-b_{-}(\bm{w}_{-},-\bm{\theta}_{-})$ and $V_{-}(\bm{w}_{-},\bm{\theta}_{-})=V_{-}(\bm{w}_{-},-\bm{\theta}_{-})$, it follows from the above discussion that we obtain
$\mathrm{MSE}_{ate}(\bm{w}_{+},\bm{w}_{-},\bm{\theta}_{+},\bm{\theta}_{-})$ $\leq$ $\mathrm{MSE}_{ate}\left(\bm{w}_{+},\bm{w}_{-},\tilde{\bm{\theta}}_{+}(\theta_{0,+}),-\bar{\bm{\theta}}_{-}(\theta_{0,-})\right)$
$=$ $\mathrm{MSE}_{ate}\left(\bm{w}_{+},\bm{w}_{-},\tilde{\bm{\theta}}_{+}(\theta_{0,+}),\tilde{\bm{\theta}}_{-}(-\theta_{0,-})\right)$
Hence, it is enough to consider the maximization of the MSE over $(\bm{\theta}_{+},\bm{\theta}_{-})\in\Theta_{+}\times\Theta_{-}$ satisfying $\theta_{0,+}\in[-1/2,0]$ and $\theta_{0,-}\in[0,1/2]$. As a result, we obtain (B.1). ∎
@@
Suppose that $\bm{w}_{+}\equiv(w_{1,+},\ldots,w_{n_{+},+})^{\prime}\in\mathcal{W}_{+}$ satisfies $w_{j,+}<w_{j+1,+}$ for some $j$. Letting $\tilde{\bm{w}}_{+}\equiv(w_{1,+},\ldots,w_{j-1,+},w_{j+1,+},w_{j,+},w_{j+2,+},\ldots,w_{n_{+},+})^{\prime}$, we have $\tilde{\bm{w}}_{+}\in\mathcal{W}_{+}$. From Lemma 2.2, we obtain

### v2

@@ Appendix A Proofs
$=$ $\min_{\bm{w}\in\mathcal{W}:\sum_{i=1}^{n}w_{i}=1}\left\{C^{2}\left(\sum_{i=1}^{n}w_{i}\|R_{i}\|^{\gamma}\right)^{2}+\frac{1}{4}\sum_{i=1}^{n}w_{i}^{2}\right\}$
$\leq$ $C^{2}\left(\frac{1}{N(\epsilon)}\sum_{i=1}^{N(\epsilon)}\|R_{i}\|^{\gamma}\right)^{2}+\frac{1}{4N(\epsilon)}$
$\leq$ $C^{2}\epsilon^{2\gamma}+\frac{1}{4N(\epsilon)}\ \leq\ C^{2}\epsilon^{2\gamma}+\frac{1}{4(c_{0}n\epsilon-n^{1-\alpha})},$
@@
and the second inequality holds since $\|R_{i}\|\leq\epsilon$ for $i=1,\dots,N(\epsilon)$. If we set $\epsilon=O(n^{-\frac{1}{2\gamma+1}})$ satisfying $\epsilon\geq 2c_{0}^{-1}n^{-\alpha}$, which exists for $\alpha\geq\frac{1}{2\gamma+1}$, then the right-hand side becomes $O(n^{-\frac{2\gamma}{2\gamma+1}})$. For example, if we set $\epsilon=2c_{0}^{-1}n^{-\frac{1}{2\gamma+1}}$, which satisfies $\epsilon\geq 2c_{0}^{-1}n^{-\alpha}$ for $\alpha\geq\frac{1}{2\gamma+1}$, then the right-hand side becomes
$4^{\gamma}C^{2}c_{0}^{-2\gamma}n^{-\frac{2\gamma}{2\gamma+1}}+\frac{1}{4(2n^{\frac{2\gamma}{2\gamma+1}}-n^{1-\alpha})}=\left(4^{\gamma}C^{2}c_{0}^{-2\gamma}+\frac{1}{4(2-n^{\frac{1}{2\gamma+1}-\alpha})}\right)n^{-\frac{2\gamma}{2\gamma+1}}=O(n^{-\frac{2\gamma}{2\gamma+1}}).$
Hence, $\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})\leq\max_{\bm{\theta}\in\Theta_{g}}\mathrm{MSE}_{g}(\tilde{\bm{w}},\bm{\theta})=O(n^{-\frac{2\gamma}{2\gamma+1}})$.
@@
where the last inequality follows from $\theta_{i}+1/2\geq 0$. This implies that $\hat{u}$ converges to one because $\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})$ converges to zero. From the finite-sample bounds (A.1), it suffices to show that
$\frac{C^{2}\sum_{i=1}^{n}\hat{w}_{i}^{2}\|R_{i}\|^{2\gamma}}{\frac{1}{4}\sum_{i=1}^{n}\hat{w}_{i}^{2}}\ \to\ 0.$ (A.2)
@@
$\frac{C^{2}\sum_{i=1}^{n}\hat{w}_{i}^{2}\|R_{i}\|^{2\gamma}}{\frac{1}{4}\sum_{i=1}^{n}\hat{w}_{i}^{2}}$ $=$ $\frac{4C^{2}\sum_{i=1}^{N(\epsilon)}\hat{w}_{i}^{2}\|R_{i}\|^{2\gamma}+4C^{2}\sum_{i=N(\epsilon)+1}^{n}\hat{w}_{i}^{2}\|R_{i}\|^{2\gamma}}{\sum_{i=1}^{n}\hat{w}_{i}^{2}}$ (A.3)
$\leq$ $4C^{2}\left\{\frac{\epsilon^{2\gamma}\left(\sum_{i=1}^{N(\epsilon)}\hat{w}_{i}^{2}\right)+\hat{w}_{N(\epsilon)}^{2}\left(\sum_{i=N(\epsilon)+1}^{n}\|R_{i}\|^{2\gamma}\right)}{\sum_{i=1}^{n}\hat{w}_{i}^{2}}\right\}$
$\leq$ $4C^{2}\left\{\epsilon^{2\gamma}+\frac{n\hat{w}_{N(\epsilon)}^{2}}{\sum_{i=1}^{n}\hat{w}_{i}^{2}}\right\},$
where the first inequality follows since $\|R_{i}\|\leq\epsilon$ for $i=1,\dots,N(\epsilon)$ and $\hat{w}_{i}\leq\hat{w}_{N(\epsilon)}$ for $i=N(\epsilon)+1,\ldots,n$ (by the extension of Lemma 2.2 for the Hölder class), and the second inequality follows since $\|R_{i}\|\leq 1$ for all $i=N(\epsilon)+1,\ldots,n$ by Assumption A.1.
@@
$O(n^{-\frac{2\gamma}{2\gamma+1}})$ $=$ $\max_{\bm{\theta}\in\Theta}\text{MSE}(\hat{\bm{w}},\bm{\theta})\ \geq\ \text{MSE}(\hat{\bm{w}},\tilde{\bm{\theta}}(0))\geq\left(\sum_{i=1}^{n}\hat{w}_{i}\min\{C\|R_{i}\|^{\gamma},1/2\}\right)^{2}$
$=$ $C^{2}\left(\sum_{i=1}^{n}\hat{w}_{i}\|R_{i}\|^{\gamma}\right)^{2}\ \geq\ C^{2}\hat{w}_{N(\epsilon)}^{2}\left(\sum_{i=1}^{N(\epsilon)}\|R_{i}\|^{\gamma}\right)^{2},$
where the second equality follows since $\hat{w}_{i}=0$ if $C\|R_{i}\|^{\gamma}\geq 1/2$ (by the extension of Lemma 2.3 for the Hölder class) and the last inequality follows since $\hat{w}_{i}\geq\hat{w}_{N(\epsilon)}$ for $i=1,\ldots,N(\epsilon)$ (by the extension of Lemma 2.2 for the Hölder class). Below, we show that
$\sum_{i=1}^{N(\epsilon)}\|R_{i}\|^{\gamma}\geq\frac{(N(\epsilon)-n^{1-\alpha})^{2}}{2c_{1}n}\ \ \text{if }\epsilon>2c_{0}^{-1}n^{-\alpha}.$ (A.5)
Once it is shown, it follows from the assumption $N(\epsilon)\geq c_{0}n\epsilon-n^{1-\alpha}$ that
$O(n^{-\frac{2\gamma}{2\gamma+1}})$ $\geq$ $C^{2}\hat{w}_{N(\epsilon)}^{2}\frac{(N(\epsilon)-n^{1-\alpha})^{4}}{4c_{1}^{2}n^{2}}$
@@
$\hat{w}_{N(\epsilon)}^{2}\leq\frac{c_{2}n^{-\frac{2\gamma}{2\gamma+1}}}{(c_{0}n^{1/2}\epsilon-2n^{1/2-\alpha})^{4}}\ \ \text{if }\epsilon>2c_{0}^{-1}n^{-\alpha}.$ (A.6)
Now we show (A.5). Since $\|R_{i}\|\leq 1$ for all $i=1,\ldots,N(\epsilon)$ by Assumption A.1, $\sum_{i=1}^{N(\epsilon)}\|R_{i}\|^{\gamma}\geq\sum_{i=1}^{N(\epsilon)}\|R_{i}\|$. Therefore, it suffices to show that, if $\epsilon>2c_{0}^{-1}n^{-\alpha}$, $\sum_{i=1}^{N(\epsilon)}\|R_{i}\|\geq\frac{(N(\epsilon)-n^{1-\alpha})^{2}}{2c_{1}n}$. Let $x^{*}(\epsilon)\equiv\frac{N(\epsilon)-n^{1-\alpha}}{c_{1}n}$, which is the unique solution to $c_{1}nx+n^{1-\alpha}=N(\epsilon)$ with respect to $x$. If $\epsilon>2c_{0}^{-1}n^{-\alpha}$, we must have $0<x^{*}(\epsilon)\leq\|R_{N(\epsilon)}\|$, since $c_{1}n\cdot 0+n^{1-\alpha}<c_{0}n\epsilon-n^{1-\alpha}\leq N(\epsilon)$ and $c_{1}n\|R_{N(\epsilon)}\|+n^{1-\alpha}\geq N(\|R_{N(\epsilon)}\|)=N(\epsilon)$ under Assumption A.1. Here, $N(\|R_{N(\epsilon)}\|)=N(\epsilon)$ holds by the definition of $N(\epsilon)$. Also, let
@@
Note that $g_{\epsilon}(x)\geq N(x)$ for all $x\in(0,\|R_{N(\epsilon)}\|]$, since $g_{\epsilon}(x)=c_{1}nx+n^{1-\alpha}\geq N(x)$ if $0<x\leq x^{*}(\epsilon)$ by Assumption A.1 and $g_{\epsilon}(x)=N(\epsilon)=N(\|R_{N(\epsilon)}\|)\geq N(x)$ if $x^{*}(\epsilon)<x\leq\|R_{N(\epsilon)}\|$. Therefore, we have
@@
Finally, combining the lower bound (A.4) on $\sum_{i=1}^{n}\hat{w}_{i}^{2}$ and the upper bound (A.6) on $\hat{w}_{N(\epsilon)}^{2}$ obtained above yields the following bound on the right-hand side of (A.3): if $\epsilon>2c_{0}^{-1}n^{-\alpha}$,
$4C^{2}\left\{\epsilon^{2\gamma}+\frac{n\hat{w}_{N(\epsilon)}^{2}}{\sum_{i=1}^{n}\hat{w}_{i}^{2}}\right\}$ $\leq$ $4C^{2}\left\{\epsilon^{2\gamma}+\frac{c_{2}n^{\frac{2\gamma+2}{2\gamma+1}}}{(c_{0}n^{1/2}\epsilon-2n^{1/2-\alpha})^{4}}\hat{u}^{-2}\right\}$
$=$ $4C^{2}\left\{\epsilon^{2\gamma}+\frac{c_{2}}{(c_{0}n^{\frac{\gamma}{4\gamma+2}}\epsilon-2n^{\frac{\gamma}{4\gamma+2}-\alpha})^{4}}\hat{u}^{-2}\right\}.$
If we set $\epsilon=n^{-\beta}$ for some $0<\beta<\min\{\frac{\gamma}{4\gamma+2},\alpha\}$ (which is equivalent to $0<\beta<\frac{\gamma}{4\gamma+2}$ as $\alpha=\frac{1}{2\gamma+1}\geq\frac{\gamma}{4\gamma+2}$), then $\epsilon>2c_{0}^{-1}n^{-\alpha}$ for any sufficiently large $n$, and we have
$4C^{2}\left\{\epsilon^{2\gamma}+\frac{c_{2}}{(c_{0}n^{\frac{\gamma}{4\gamma+2}}\epsilon-2n^{\frac{\gamma}{4\gamma+2}-\alpha})^{4}}\hat{u}^{-2}\right\}$ $=$ $4C^{2}\left\{n^{-2\beta\gamma}+\frac{c_{2}}{(c_{0}n^{\frac{\gamma}{4\gamma+2}-\beta}-2n^{\frac{\gamma}{4\gamma+2}-\alpha})^{4}}\hat{u}^{-2}\right\}$
$=$ $o(1).$
@@
where $\bm{Y}$ and $\tilde{\bm{Y}}$ follow Bernoulli distribution with parameters $\bm{p}$ and $\tilde{\bm{p}}(p_{0,+},\tau_{0})$, respectively. Then, $\hat{\tau}(\bm{Y})$ is increasing function in $Y_{i,+}$ and decreasing in $Y_{i,-}$. Because $\tilde{Y}_{i,+}$ has first-order stochastic dominance over $Y_{i,+}$ and $\tilde{Y}_{i,-}$ has first-order stochastic dominance over $Y_{i,-}$, it follows from Lemma 1 of ishihara2023bandwidth that we have
@@ Appendix B Minimax estimation for the average treatment effect
Fix $\theta_{0,+}\in\Theta_{+}$ and $\theta_{0,-}\in\Theta_{-}$. Without loss of generality, we assume $\theta_{0,+}\leq 0$. Define
$\bar{\bm{\theta}}_{+}(\theta_{0,+})$ $\equiv$ $\left(\theta_{0,+},\max\{\theta_{0,+}-C\|R_{1,+}\|,-1/2\},\ldots,\max\{\theta_{0,+}-C\|R_{n_{+},+}\|,-1/2\}\right)^{\prime},$
$\bar{\bm{\theta}}_{-}(\theta_{0,-})$ $\equiv$ $\left(\theta_{0,-},\min\{\theta_{0,-}+C\|R_{1,-}\|,1/2\},\ldots,\min\{\theta_{0,-}+C\|R_{n_{-},-}\|,1/2\}\right)^{\prime}.$
First, we consider the case where $\theta_{0,+}\in[-1/2,0]$ and $\theta_{0,-}\in[0,1/2]$. In this case, Theorem 2.1 implies that $b_{+}(\bm{w}_{+},\bm{\theta}_{+})^{2}+V_{+}(\bm{w}_{+},\bm{\theta}_{+})$ is maximized at $\bm{\theta}_{+}=\tilde{\bm{\theta}}_{+}(\theta_{0,+})$ and $b_{-}(\bm{w}_{-},\bm{\theta}_{-})^{2}+V_{-}(\bm{w}_{-},\bm{\theta}_{-})$ is maximized at $\bm{\theta}_{-}=\tilde{\bm{\theta}}_{-}(\theta_{0,-})$. Note that
$b_{+}(\bm{w}_{+},\bm{\theta}_{+})=\sum_{i=1}^{n_{+}}w_{i,+}(\theta_{i,+}-\theta_{0,+})-\left(1-\sum_{i=1}^{n_{+}}w_{i,+}\right)\theta_{0,+}$
and
$b_{-}(\bm{w}_{-},\bm{\theta}_{-})=\sum_{i=1}^{n_{-}}w_{i,-}(\theta_{i,-}-\theta_{0,-})-\left(1-\sum_{i=1}^{n_{-}}w_{i,-}\right)\theta_{0,-}.$
Then $|b_{+}(\bm{w}_{+},\bm{\theta}_{+})|$ is maximized at $\bm{\theta}_{+}=\tilde{\bm{\theta}}_{+}(\theta_{0,+})$ or $\bar{\bm{\theta}}_{+}(\theta_{0,+})$ and $|b_{-}(\bm{w}_{-},\bm{\theta}_{-})|$ is maximized at $\bm{\theta}_{-}=\tilde{\bm{\theta}}_{-}(\theta_{0,-})$ or $\bar{\bm{\theta}}_{-}(\theta_{0,-})$. Because $\bm{w}_{+}\in\mathcal{W}_{+}$ and $\theta_{0,+}\leq 0$, we have $|b_{+}(\bm{w}_{+},\bar{\bm{\theta}}_{+}(\theta_{0,+}))|\leq|b_{+}(\bm{w}_{+},\tilde{\bm{\theta}}_{+}(\theta_{0,+}))|=b_{+}(\bm{w}_{+},\tilde{\bm{\theta}}_{+}(\theta_{0,+}))$. Similarly, we have $|b_{-}(\bm{w}_{-},\bar{\bm{\theta}}_{-}(\theta_{0,-}))|\leq|b_{-}(\bm{w}_{-},\tilde{\bm{\theta}}_{-}(\theta_{0,-}))|=-b_{-}(\bm{w}_{-},\tilde{\bm{\theta}}_{-}(\theta_{0,-}))$. These results imply that $b_{+}(\bm{w}_{+},\bm{\theta}_{+})b_{-}(\bm{w}_{-},\bm{\theta}_{-})$ is maximized at $\bm{\theta}_{+}=\tilde{\bm{\theta}}_{+}(\theta_{0,+})$ and $\bm{\theta}_{-}=\tilde{\bm{\theta}}_{-}(\theta_{0,-})$. Therefore, if $\theta_{0,+}\in[-1/2,0]$ and $\theta_{0,-}\in[0,1/2]$ hold, we obtain
@@
Moreover, since $\theta_{0,+}\leq 0$ and $\theta_{0,-}\leq 0$, it follows from an argument similar to the one above that $|b_{+}(\bm{w}_{+},\bm{\theta}_{+})|\leq|b_{+}(\bm{w}_{+},\tilde{\bm{\theta}}_{+}(\theta_{0,+}))|=b_{+}(\bm{w}_{+},\tilde{\bm{\theta}}_{+}(\theta_{0,+}))$ and $|b_{-}(\bm{w}_{-},\bm{\theta}_{-})|\leq|b_{-}(\bm{w}_{-},\bar{\bm{\theta}}_{-}(\theta_{0,-}))|=b_{-}(\bm{w}_{-},\bar{\bm{\theta}}_{-}(\theta_{0,-}))$. Hence,
$\mathrm{MSE}_{ate}(\bm{w}_{+},\bm{w}_{-},\bm{\theta}_{+},\bm{\theta}_{-})$
$\hskip 14.45377pt+b_{-}(\bm{w}_{-},\bar{\bm{\theta}}_{-}(\theta_{0,-}))^{2}+V_{-}(\bm{w}_{-},\bar{\bm{\theta}}_{-}(\theta_{0,-}))+2b_{+}(\bm{w}_{+},\tilde{\bm{\theta}}_{+}(\theta_{0,+}))b_{-}(\bm{w}_{-},\bar{\bm{\theta}}_{-}(\theta_{0,-})).$
Because $b_{-}(\bm{w}_{-},\bm{\theta}_{-})=-b_{-}(\bm{w}_{-},-\bm{\theta}_{-})$, $V_{-}(\bm{w}_{-},\bm{\theta}_{-})=V_{-}(\bm{w}_{-},-\bm{\theta}_{-})$, and $\bar{\bm{\theta}}_{-}(\theta_{0,-})=\tilde{\bm{\theta}}_{-}(-\theta_{0,-})$, we obtain
$\mathrm{MSE}_{ate}(\bm{w}_{+},\bm{w}_{-},\bm{\theta}_{+},\bm{\theta}_{-})$
$\leq$ $b_{+}(\bm{w}_{+},\tilde{\bm{\theta}}_{+}(\theta_{0,+}))^{2}+V_{+}(\bm{w}_{+},\tilde{\bm{\theta}}_{+}(\theta_{0,+}))+b_{-}(\bm{w}_{-},-\bar{\bm{\theta}}_{-}(\theta_{0,-}))^{2}+V_{-}(\bm{w}_{-},-\bar{\bm{\theta}}_{-}(\theta_{0,-}))$
$\hskip 14.45377pt-2b_{+}(\bm{w}_{+},\tilde{\bm{\theta}}_{+}(\theta_{0,+}))b_{-}(\bm{w}_{-},-\bar{\bm{\theta}}_{-}(\theta_{0,-}))$
$=$ $\mathrm{MSE}_{ate}\left(\bm{w}_{+},\bm{w}_{-},\tilde{\bm{\theta}}_{+}(\theta_{0,+}),-\bar{\bm{\theta}}_{-}(\theta_{0,-})\right)$
$=$ $\mathrm{MSE}_{ate}\left(\bm{w}_{+},\bm{w}_{-},\tilde{\bm{\theta}}_{+}(\theta_{0,+}),\tilde{\bm{\theta}}_{-}(-\theta_{0,-})\right).$
Hence, it is enough to consider the maximization of $\mathrm{MSE}_{ate}\left(\bm{w}_{+},\bm{w}_{-},\tilde{\bm{\theta}}_{+}(\theta_{0,+}),\tilde{\bm{\theta}}_{-}(\theta_{0,-})\right)$ over $\theta_{0,+}\in[-1/2,0]$ and $\theta_{0,-}\in[0,1/2]$. As a result, we obtain (B.1). ∎
@@
By Theorem B.1, it suffices to consider the MSE at $(\tilde{\bm{\theta}}_{+}(\theta_{0,+}),\tilde{\bm{\theta}}_{-}(\theta_{0,-}))$ for $\theta_{0,+}\in[-1/2,0]$ and $\theta_{0,-}\in[0,1/2]$. Suppose that $\bm{w}_{+}\equiv(w_{1,+},\ldots,w_{n_{+},+})^{\prime}\in\mathcal{W}_{+}$ satisfies $w_{j,+}<w_{j+1,+}$ for some $j$. Letting $\tilde{\bm{w}}_{+}\equiv(w_{1,+},\ldots,w_{j-1,+},w_{j+1,+},w_{j,+},w_{j+2,+},\ldots,w_{n_{+},+})^{\prime}$, we have $\tilde{\bm{w}}_{+}\in\mathcal{W}_{+}$. From the proof of Lemma 2.2, for any $\theta_{0,+}\in[-1/2,0]$,